VoIP Disruption

Posted: August 23rd, 2017 at 10:14 by Adam Heath

We are aware of an issue currently affecting some VoIP services.  We are investigating the issue and will provide further information once it is available.

  1. Adam Heath says:

    Our investigations have led to an issue being identified with a specific upstream carrier. Our engineers are in correspondence with the relevant carrier for more information, which we will make available once obtained.

  2. Adam Heath says:

    Further information attained suggests the root cause may be a power issue with an upstream carrier. It has been advised that there are engineers working to restore service.

  3. Adam Heath says:

    The affected upstream carrier has confirmed that a small number of services have been restored, but there is still a significant amount of non-functional incoming services. Work is continuing towards restoration.

  4. Adam Heath says:

    We have started to see inbound services recover, following a way to reroute numbers in bulk being identified further upstream from ourselves. Internal testing here against numbers previously reported to us, also confirms this to be the case.

  5. Adam Heath says:

    An incident report for this issue is now available here.

At-Risk: Kingston to Slough fibre link maintenance

Posted: August 16th, 2017 at 07:38 by Iain Beveridge

21-Aug-2017 02:00 GMT to 21-Aug-2017 06:00 GMT

During the above window one of our fibre suppliers will be carrying out maintenance on one of our links between Kingston and Slough. Whilst the actual downtime is expected to be less, the full 4 hrs may be required to account for any unforeseen issues. This work is not expected to be service affecting, however, you may see unusual paths and increase in latency while traffic is automatically rerouted via alternate routes.

Incident: Interxion

Posted: July 28th, 2017 at 09:42 by Richard Partridge

We are currently experiencing a loss of connectivity to one of our switches in Interxion. Engineers are already investigating and local engineering resources are on the way to site. Further updates will be provided as and when they become available.

  1. Richard Partridge says:

    Connectivity has been restored, however we are still investigating the underlying cause.

Incident: VoIP Disruption

Posted: July 13th, 2017 at 10:46 by Richard Partridge

Engineers are currently investigating an incident that is causing a loss of service on our VoIP platform. Further information will be provided as and when it becomes available.

  1. Richard Partridge says:

    Service has now been restored. We apologise for any inconvenience.

Email Disruption

Posted: July 13th, 2017 at 10:15 by Richard Partridge

We are currently receiving reports that a small number of DSL customers are experiencing difficulty when trying to send or receive emails. Our engineers are currently investigating and further information will be provided as and when it becomes available.

  1. Richard Partridge says:

    Engineers are still investigating the email issue at this time. Further updates will be provided over the course of the day.

  2. Richard Partridge says:

    Engineers are continuing their investigation. Further updates will be provided over the course of the day

  3. Richard Partridge says:

    Engineers have now been able to recreate the symptoms being reported and are investigating a possible root cause of the problem. Further updates will follow in due course.

  4. Richard Partridge says:

    Adjustments have been made to try and resolve this problem, however we will need to return to normal bandwidth levels in order to determine their effectiveness. Further updates will be provided tomorrow morning once the impact of these changes have been ascertained.

  5. Richard Partridge says:

    “Yesterday’s adjustments have had a positive affect, but haven’t entirely resolved the problem. We’re continuing to work on this and will provide further updates over the course of the day.”

  6. Richard Partridge says:

    No further issues have been reported over the course of the afternoon. We will closely monitor the situation over the course of the weekend and advise of any further occurrences of this issue accordingly.

BT circuit issue

Posted: July 6th, 2017 at 10:56 by Adam Heath

We are aware of an issue affecting a number of BT Ethernet circuits which connect to our Columbo core.   Our initial diagnostics indicate the issue to be on BT’s hub and we have this raised with them already for investigation.

We will provide further information once it becomes available.

  1. Adam Heath says:

    We are continuing to correspond and troubleshoot with the suppliers. As a precaution, we have also tasked out an engineer to double check our equipment at the site.

    Further updates will be available shortly.

  2. Adam Heath says:

    We have restored service to affected circuits following an engineer replacing a faulty optic at site. We’re continuing to monitor from our side.

Planned Maintenance: Manchester

Posted: June 28th, 2017 at 16:19 by Mark Yardley

Thursday 6th July 2017 00:01 – 06:00

During the above window we will be replacing man.core.enta.net with a higher capacity router.

All services currently terminated in this location will be migrated on to the new hardware. This includes all customer transit, EWCS, wholesale hand-off and pseudowires terminating at this location.

All other traffic that would normally route through this location will take alternative paths.

  1. Richard Partridge says:

    This work has been postponed and will be rescheduled for a later date.

Incident: Viper Hosting Platform

Posted: June 26th, 2017 at 11:53 by Adam Heath

We are currently investigating an issue affecting some of our Viper hosting platform.

Engineers are currently investigating and further updates will be provided when available.

  1. Adam Heath says:

    Our NOC engineers have started to restore service to the Viper platform which requires a reboot of each server. Current focus is on returning the functionality for users, before reviewing to confirm the root cause.

    Further information will be available on this shortly.

  2. Adam Heath says:

    Our NOC engineers have confirmed that almost all users should now have functionality restored, with the final few affected servers still in the process of being rebooted and brought back online. This is expected to be complete within the next hour.

    Once complete, we’ll be ascertaining the cause and providing the information via this site.

Planned Maintenance: Faraday.core

Posted: June 26th, 2017 at 09:03 by Iain Beveridge

Wednesday 05 July 2017 00:00 – 4:00

During the above window we will be moving some customers with EAD circuits directly connected to faraday.core from one switch to another. These customers will be contacted directly before the window and should only experience a small amount of downtime each. However as a precaution we would consider all customers terminating in Faraday at risk.

  1. This work has now been completed and passed the post maintenance check list to ensure customer connectivity has been restored to effected customers.

Incident: Bristol.core reload

Posted: June 21st, 2017 at 23:02 by Neil Watson

We are aware that our Bristol.core router has unexpectedly reloaded. Customers routed via this node will have temporarily lost service while the reload completed. Currently, we do not believe that the reload is related to the ongoing planned work. We will investigate the cause and will report as appropriate. Apologies for the inconvenience caused.